### v1

# Measurement-aligned Flow for Inverse Problem

Abstract
Diffusion models provide a powerful way to incorporate complex prior information for solving inverse problems. However, existing methods struggle to correctly incorporate guidance from conflicting signals in the prior and measurement, especially in the challenging setting of non-Gaussian or unknown noise. To bridge these gaps, we propose Measurement-Aligned Sampling (MAS), a novel framework for linear inverse problem solving that can more flexibly balance prior and measurement information. MAS unifies and extends existing approaches like DDNM and DAPS, and offers a new optimization perspective. MAS can generalize to handle known Gaussian noise, unknown or non-Gaussian noise types. Extensive experiments show that MAS consistently outperforms state-of-the-art methods across a range of tasks.
1 Introduction
Inverse problems are prevalent in image restoration (IR) tasks, including super-resolution, inpainting, deblurring, colorization, denoising, and JPEG restoration (chung2022diffusion, ; kawar2022denoising, ; saharia2022image, ; wang2022zero, ; lugmayr2022repaint, ; mardani2023variational, ; song2023pseudoinverse, ; kawar2022jpeg, ). Solving an inverse problem involves recovering an unknown original image $x_{0}\in\mathbb{R}^{n}$ based on information from a prior distribution, $\pi(x_{0})$, and noisy measurements $y\in\mathbb{R}^{m}$ generated through a forward model:
$y=\mathcal{H}(x_{0})+\epsilon.$ (1)
Here ${\epsilon\in\mathbb{R}^{m}}$ represents measurement noise, $x_{0}\in\mathbb{R}^{d}$ is drawn from data distribution $\pi_{0}(x_{0})$, $\mathcal{H}:\mathbb{R}^{d}\mapsto\mathbb{R}^{m}$ is the measurement function, and $y\in\mathbb{R}^{m}$ denotes the degraded measurement or observed image. A useful motivating example is a high-resolution image $x_{0}$, with a noisy degraded image $y$ and known corruption process.
Pretrained diffusion or flow models provide a prior distribution $\pi_{0}(x_{0})$ that significantly facilitates the solution of inverse problems. DPS (chung2022diffusion, ) and $\Pi$GDM (song2023pseudoinverse, ) estimate conditional scores directly from the measurement model by leveraging score decomposition to guide each diffusion sampling iteration. In contrast, FPS (dou2024diffusion, ), DAPS (zhang2024improving, ), and MPGD (he2023manifold, ) align denoiser outputs directly with measurements, thereby avoiding backpropagation. However, these methods are not explicitly designed to handle high levels of Gaussian noise, as addressed by approaches like DDNM (wang2022zero, ) and DDRM (kawar2022denoising, ).
Moreover, the above approaches lack the ability to effectively regulate the ‘strength’ and ‘type’ of the alignment between measurements and model predictions. They also commonly assume a known forward operator $\mathcal{H}$ and Gaussian noise with known variance, i.e., $\epsilon\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I}_{m})$. In practical settings, noise frequently deviates from Gaussian assumptions—exhibiting characteristics like salt-and-pepper, periodic, or Poisson distributions—or is completely unknown. Additionally, the forward measurement operator may be uncertain or inaccurately specified. Effectively addressing inverse problems under these more general and realistic conditions remains an open and challenging research area.
Our main contributions are summarized as follows:
  • We propose Measurement-Aligned Sampling (MAS), a novel framework for solving linear inverse problems. MAS provides additional optimization perspectives and generalizes approaches such as DDNM and DAPS for linear inverse problems. Furthermore, our proposed “overshooting” technique achieves superior restoration quality compared to DDNM across various inverse problem scenarios.
  • We develop new techniques to handle Gaussian noise and unknown noise. Specifically, we extend strategies akin to DDNM$+$ to address noisy inverse problems with known Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{y}\mathbb{I}_{m})$. Moreover, our novel parameterization scheme allows us to effectively handle noisy inverse problems with unknown or non-Gaussian noise structures and even non-differentiable measurements, such as JPEG restoration, without requiring explicit knowledge of the forward operator or noise level. The comparison of method applicability across different inverse problems is shown in Table 1.
  • Our experimental results demonstrate state-of-the-art performance across a diverse set of tasks, see examples in Fig. 1 and experiments in Sec. 5.
Figure 1: Solving various inverse problems using unconditional diffusion models. Our model demonstrates better robustness with unknown noise and strong Gaussian noise.
Table 1: Comparison of method applicability across different inverse problems.
Inverse Problem Noise strength DDNM DDRM $\Pi$GDM DAPS RED-Diff MAS (ours)
Linear + noise free -
Linear + Gaussian noise Known
Linear + non-Gaussian noise Unknown
JPEG / Quantization restoration Known
JPEG / Quantization restoration Unknown
2 Background
Notation. Our time notation follows diffusion literature where $t=T$ corresponds to the source distribution and $t=0$ corresponds to the target distribution, consistent with the use of pretrained diffusion models in our experiments. In contrast, flow-based models typically adopt the reverse, with $t=0$ as the source distribution.
2.1 Conditional Diffusion and flow models
Given training dataset $\mathcal{D}=\{x_{0}^{i}\}_{i=1}^{N}$ from target distribution $\pi_{0}(x_{0})$, $x_{0}^{i}\in\mathbb{R}^{d}$, the goal of generative modeling is to draw new samples from $\pi_{0}$. In the context of conditional generation, suppose that we have data samples from a joint disribution $(x_{0}^{i},y)\sim\pi(x_{0},y)$, where $x_{0}$, $y$ are dependent, and $y$ could be class labels or text information, for example.
[…]
$dx=\left[(\dot{\alpha}_{t}-\alpha_{t}\xi_{t})\tilde{x}_{0}(t,x,y)+\xi_{t}x_{t}\right]dt+\epsilon_{t}dW_{t},$ (2)
where $\tilde{x}_{0}(t,x_{t},y)=\mathbb{E}[x_{0}\mid x_{t},y]$ and $\xi_{t}=\frac{\dot{\sigma}_{t}}{\sigma_{t}}-\frac{\epsilon_{t}^{2}}{2\sigma_{t}^{2}}$. Then $x_{0}$ could be sampled by solving the SDE described in Eq. 2. Although different sampling methods can be chosen, generally, the iteration follows the form:
$x_{t-\Delta t}\sim\mathcal{N}(a_{t}\tilde{x}_{0}(t,x,y)+b_{t}x_{t},c_{t}^{2}\mathbb{I}).$ (3)
2.2 Training-free conditional inference
Given pretrained unconditional diffusion models $\hat{x}_{0}^{\theta}\approx\mathbb{E}[x_{0}\mid x_{t}]$, training-free approaches for conditional inference allows us to approximate the ideal denoiser $\tilde{x}_{0}(t,x_{t},y)$ (pokle2023training, ):
$\tilde{x}_{0}(t,x_{t},y)=\mathbb{E}[x_{0}\mid x_{t},y]\approx\hat{x}_{0}^{\theta}+\frac{\sigma_{t}^{2}}{\alpha_{t}}\nabla_{x_{t}}\log p(y\mid x_{t}).$ (4)
For inverse problems or conditioning problems where $p(y\mid x_{0})$ is known but $p(y\mid x_{t})$ is intractable (chung2022diffusion, ; wu2024practical, ), denoising predictors $\hat{x}_{0}^{\theta}(t,x_{t})$ provide a convenient single-sample approximation for $p(y|x_{t})$, $p(y\mid x_{t})\approx p(y\mid\hat{x}_{0}^{\theta}(t,x_{t}))$. For example, $\Pi$GDM (song2023pseudoinverse, ) assumes $p_{t}(x_{0}\mid x_{t})\approx\mathcal{N}(\hat{x}_{0}^{\theta}(t,x_{t}),r_{t}^{2}\mathbb{I})$, then
[…]
Then the gradient guidance can be approximated by differentiating through the denoising predictor network using the chain rule:
$\nabla_{x_{t}}\log p(y\mid x_{t})\approx\left[(y-\hat{x}_{0}^{\theta})^{T}(r_{t}^{2}HH^{\mathsf{T}}+\sigma_{y}^{2}\mathbb{I})H\frac{\partial\hat{x}_{0}^{\theta}(t,x_{t})}{\partial x_{t}}\right]^{\mathsf{T}}.$ (6)
Substituting Eq. 6 into Eq. 4, we obtain an estimate of the ideal denoiser $\tilde{x}_{0}(t,x_{t},y)$. While this method works well in practice to produce the desired guidance effect (Sec. 5), it still relies on a crude approximation of the expectation in Eq. 5 and requires backpropagation through the denoising network. In this work, we present an alternative approach to estimate $\tilde{x}_{0}(t,x_{t},y)$ based on unconditional denoising predictors $\hat{x}_{0}^{\theta}(t,x_{t})$ and the measurement $y$.
3 Methodology
3.1 Measurement-Aligned Sampling for linear inverse problem
Measurement-Aligned Sampling aims to estimate the desired conditional denoiser $\tilde{x}_{0}(t,x_{t},y)=\mathbb{E}[x_{0}\mid x_{t},y]$ at time $t$ by utilizing two measurements: an approximated unconditional denoiser $\hat{x}^{\theta}_{0}(t,x_{t})\approx\mathbb{E}[x_{0}\mid x_{t}]$ and the measurement $y$. MAS performs posterior sampling $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$ based on the following probabilistic model:
Prior: $\quad p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})\coloneqq\mathcal{N}(\hat{x}_{0}^{\theta},\rho_{t}^{2}\mathbb{I}_{d}),$ (7a)
[…]
where $\rho_{t}$ is a hyperparameter defining the prior variance.
Typically, computing the exact likelihood in Eq. 7b is infeasible since $p(x_{0}\mid\tilde{x}_{0})$ is unknown. $\Pi$GDM (song2023pseudoinverse, ) assume $p(x_{0}\mid\hat{x}_{0}^{\theta})=\mathcal{N}(\hat{x}_{0}^{\theta},r_{t}^{2}\mathbb{I})$, where $\hat{x}_{0}^{\theta}\approx\mathbb{E}[x_{0}\mid x_{t}]$. In our work, we revise this assumption to $p(x_{0}\mid\tilde{x}_{0})=\mathcal{N}(\tilde{x}_{0},r_{t}^{2}\mathbb{I})$. Consequently, for linear inverse problem where $p(y\mid x_{0})\coloneqq\mathcal{N}(Hx_{0},\sigma_{y}^{2}\mathbb{I})$, the likelihood $p(y\mid\tilde{x}_{0})$ has similar Gaussian form as in Eq. 5:
$p(y\mid\tilde{x}_{0})=\mathcal{N}(y;H\tilde{x}_{0},\sigma_{y}^{2}\mathbb{I}_{m}+r_{t}^{2}HH^{\mathsf{T}}).$ (8)
Given the prior $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})$ in Eq. 7a and the likelihood in Eq. 7b, the posterior distribution $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$ can be analytically derived via Bayesian linear regression, as detailed in Prop. 3.1. See App. A for the full derivation and proofs.
Proposition 3.1 (Bayesian Linear Regression).
If $p(y\mid\tilde{x}_{0})$ and $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})$ are both Gaussian or can be approximated as such, $p({y}\mid\tilde{x}_{0})=\mathcal{N}(H\tilde{x}_{0},\sigma_{y}^{2}{\mathbb{I}}+r_{t}^{2}HH^{\mathsf{T}})$, $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})=\mathcal{N}(\hat{x}_{0}^{\theta},\rho_{t}^{2}{\mathbb{I}_{d}})$, then $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)=\mathcal{N}(\mu_{\mathrm{post}},\Sigma_{\mathrm{post}})$:
$\Sigma_{\mathrm{post}}$ $=\left(\frac{1}{\rho_{t}^{2}}\mathbb{I}_{d}+H^{\mathsf{T}}(r_{t}^{2}HH^{\mathsf{T}}+\sigma_{y}^{2}\mathbb{I}_{m})^{-1}H\right)^{-1},$ (9)
$\mu_{\mathrm{post}}$ $=\Sigma_{\mathrm{post}}\left[\frac{1}{\rho_{t}^{2}}\hat{x}_{0}^{\theta}+H^{\mathsf{T}}(r_{t}^{2}HH^{\mathsf{T}}+\sigma_{y}^{2}\mathbb{I}_{m})^{-1}y\right].$ (10)
This result provides a concrete way to sample $\tilde{x}_{0}$ from the posterior $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})$ without backpropagation through the denoising network, as is needed in $\Pi$GDM (song2023pseudoinverse, ). We discuss implementation of this scheme before generalizing to non-Gaussian noise models in Sec. 4.
3.2 Implementation
Posterior sampling.
Given the posterior distribution $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$ explicitly derived via Bayesian linear regression in Prop. 3.1, we can efficiently sample $\tilde{x}_{0}\sim p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$. Defining $\eta_{1}\coloneqq{r_{t}^{2}}/{\rho_{t}^{2}}$, $\eta_{2}\coloneqq\sigma_{y}^{2}/\rho_{t}^{2}$, posterior sampling $\tilde{x}_{0}^{\text{ps}}\sim\mathcal{N}(\mu_{\text{post}},\Sigma_{\text{post}})$ can be implemented by a single draw of $\epsilon_{p}\sim\mathcal{N}(0,\mathbb{I}_{d})$:
$\tilde{x}_{0}^{\text{ps}}$ $=Y^{-1}[\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y+\rho_{t}\epsilon_{p}],\quad\epsilon_{p}\sim\mathcal{N}(0,\mathbb{I}_{d})$ (11)
$\text{where}\quad W$ $\coloneqq\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m},\quad Y\coloneqq\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H,$
then $\tilde{x}_{0}^{\text{ps}}\sim p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y)$.
Posterior mean.
In the implementation, posterior sampling in Eq. 11 introduces additional noise $\rho_{t}Y^{-1}\epsilon_{p}$, which leads to noisy restoration in the experiment, See App. G for more details. For simplicity and improved reconstruction quality, we could directly replace $\tilde{x}_{0}^{\text{ps}}$ in Eq. 11 by the posterior mean $\tilde{x}_{0}^{\text{pm}}\coloneqq\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]$:
$\tilde{x}_{0}^{\text{pm}}=Y^{-1}[\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (12)
Efficient linear solves via the SVD of $H$.
Computing the inverse $W^{-1}$ and $Y^{-1}$ in Eq. 11 and Eq. 12 naively can be computationally expensive. Thus, in practice, we would like to avoid directly computing the inverse. Let $H=U\Sigma V^{\mathsf{T}}$ with singular values $s_{1},\dots,s_{n}$. Then
[…]
see App. A for the proof. Hence both $W^{-1}v$ and $Y^{-1}u$ reduce to inexpensive diagonal scalings in the SVD basis, avoiding the calculation of any explicit matrix inversion or square-root.
3.3 Exploring alignment trade-offs via an optimization perspective
We can interpret the posterior mean $\tilde{x}_{0}^{\text{pm}}$ from an optimization perspective to explicitly see how our approach incorporates alignment from the prior and the measurement. Specifically, $\tilde{x}_{0}^{\text{pm}}$ minimizes the following objective function (see App. A for the proof):
$\tilde{x}_{0}^{\text{pm}}=\arg\min\{\|\tilde{x}_{0}-\hat{x}_{0}^{\theta}\|^{2}+\|y-H\tilde{x}_{0}\|_{W^{-1}}^{2}\}$ (14)
where $W=\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m}$ (with $\eta_{1}\geq 0,\eta_{2}\geq 0$) emerges as a metric that reflects the relative importance of measurement fidelity and prior regularization. As $\eta_{1}=0,\eta_{2}>0$, Eq. 14 reduced to:
$\tilde{x}_{0}^{\text{pm}}=\arg\min\{\|\tilde{x}_{0}-\hat{x}_{0}^{\theta}\|^{2}+\frac{1}{\eta_{2}}\|y-H\tilde{x}_{0}\|^{2}\}$ (15)
This reduced form represents a standard Tikhonov regularization (ridge regression) problem. The choice of $\eta_{2}$ here controls the trade-off between fidelity to the measurements $y$ and $\hat{x}_{0}^{\theta}$. While as $\eta_{1}>0,\eta_{2}=0$, the optimization takes on the following generalized least squares (GLS) form:
$\tilde{x}_{0}^{\text{pm}}=\arg\min\{\|\tilde{x}_{0}-\hat{x}_{0}^{\theta}\|^{2}+\frac{1}{\eta_{1}}\|y-H\tilde{x}_{0}\|_{(HH^{\mathsf{T}})^{-1}}^{2}\}$ (16)
Here, the weighting term $(HH^{\mathsf{T}})^{-1}$ indicates a Mahalanobis-type distance metric, emphasizing directions corresponding to smaller singular values of $H$. The choice of $\eta_{1}$ and $\eta_{2}$ significantly affects the reconstruction quality, as will be demonstrated experimentally.
Remark 1 (Connection with DAPS (zhang2024improving, )).
As $\eta_{1}=0$, $\tilde{x}_{0}^{\text{ps}}$ is reduced to posterior sampling in DAPS.
Remark 2 (Connection with DDNM (wang2022zero, )).
As $\eta_{2}=0$ and $\eta_{1}\rightarrow 0$, $\tilde{x}_{0}^{\text{pm}}\rightarrow\tilde{x}_{0}^{\text{DDNM}}=\hat{x}_{0}^{\theta}+H^{\dagger}(y-H\hat{x}_{0}^{\theta})$. Thus, in this limiting case, MAS recovers DDNM.
Figure 2: 2D illustration of the influence of parameters $\eta_{1}$ and $\eta_{2}$. Dots represent posterior means $\tilde{x}_{0}^{\text{pm}}$, calculated via Eq. 12. (a) Parameter $\eta_{1}$ controls the trade-off between $\hat{x}_{0}^{\theta}$ and $\tilde{x}_{0}^{\text{DDNM}}$: as $\eta_{1}\rightarrow\infty$, the posterior mean $\tilde{x}_{0}^{\text{pm}}$ approaches $\hat{x}_{0}^{\theta}$; as $\eta_{1}\rightarrow 0$, it converges to $\tilde{x}_{0}^{\text{DDNM}}$. (b) Adjusting $\eta_{2}$ alters the posterior trajectory differently from varying $\eta_{1}$.
Remark 3 (Overshooting trick).
Theoretically, $\eta_{1}\geq 0$ since $\eta_{1}\coloneqq r_{t}^{2}/\rho_{t}^{2}$, however, the posterior mean in Eq. 12 allows negative $\eta_{1}$. As shown in Fig. 2, negative $\eta_{1}$ leads to overshooting, which demonstrates better performance than DDNM in the experiments.
4 Generalizing to noisy inverse problems
4.1 Identifying the introduced noise
To illustrate our ideas, we first consider a simple forward model: $y=x_{0}+\epsilon_{y}$. Calculating the posterior mean $\tilde{x}_{0}^{\text{pm}}$ (Eq. 12) leads to:
$\tilde{x}_{0}^{\text{pm}}=\hat{x}_{0}^{\theta}+\frac{y-\hat{x}_{0}^{\theta}}{\eta_{1}+\eta_{2}+1}=\hat{x}_{0}^{\theta}+\frac{x_{0}-\hat{x}_{0}^{\theta}}{\eta_{1}+\eta_{2}+1}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\epsilon_{y}}{\eta_{1}+\eta_{2}+1}},$ (17)
where $\epsilon_{y}/(\eta_{1}+\eta_{2}+1)$ is the additional noise introduced to $\tilde{x}_{0}^{\text{pm}}$, and will be further introduced into $x_{t-\Delta t}$. To address this issue, we propose two approaches for addressing known Gaussian noise (Sec. 4.2) and unknown noise Sec. 4.3.
4.2 Addressing Gaussian noise with known variance
To handle Gaussian noise with known variance, we modify Eq. 17 and Eq. 3 as:
$\tilde{x}_{0}=\hat{x}_{0}^{\theta}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda_{t}}\frac{y-\hat{x}_{0}^{\theta}}{\eta_{1}+\eta_{2}+1},\quad x_{t-\Delta t}\sim\mathcal{N}(a_{t}\tilde{x}_{0}+b_{t}x_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t}}\mathbb{I}).$ (18)
Here $\lambda_{t}$ and $r_{t}$ are two parameters that can control the total noise introduced to $x_{t-\Delta t}$. In our work, we adopt similar two principles as DDNM$+$ (wang2022zero, ) for handling Gaussian noise: (i) the total noise introduced in $x_{t-\Delta t}$ should be $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$ to conform the correct distribution of $x_{t-\Delta t}$ in Eq. 3; (ii) $\lambda_{t}$ should be as close to $1$ as possible to maximize the preservation of $\tilde{x}_{0}^{\text{pm}}$. As $\epsilon_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I})$, principle (i) and principle (ii) are equivalent to:
$\left(\frac{a_{t}\lambda_{t}\sigma_{y}}{\eta_{1}+\eta_{2}+1}\right)^{2}+\gamma_{t}$ $=c_{t}^{2},\quad\lambda_{t}=\begin{cases}1,&c_{t}\geq\dfrac{a_{t}\sigma_{y}}{\eta_{1}+\eta_{2}+1}\\[4.0pt] \dfrac{c_{t}(\eta_{1}+\eta_{2}+1)}{a_{t}\sigma_{y}},&c_{t}<\dfrac{a_{t}\sigma_{y}}{\eta_{1}+\eta_{2}+1}\end{cases}.$ (19)
The more accurate version for general form of $H$ could be found in App. B. Note that the revision didn’t introduce additional parameters.
4.3 Addressing unknown noise and non-differentiable measurements
Figure 3: The sample process of solving inverse problem with unknown noise. Here we set $\eta_{1}=0$ and $\eta_{2}=0.5a_{t}/c_{t}$.
Addressing unknown noise or non-Gaussian noise. When the measurement noise $\sigma_{y}$ is non-Gaussian or unknown, it becomes difficult to ensure that the total noise in $x_{t-\Delta t}$ follows the desired distribution $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$, To address this, we continue to sample $x_{t-\Delta t}$ using Eq. 3. Next, the noise introduced to $x_{t-\Delta t}$ is the sum of two components:
$\epsilon_{\text{ng}}=(a_{t}\lambda_{t}\epsilon_{y})/(\eta_{1}+\eta_{2}+1),\quad\epsilon_{\text{g}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I}).$ (20)
Here $\epsilon_{\text{ng}}$ is related to the noise introduced by unknown noise $\epsilon_{y}$, while $\epsilon_{\text{g}}$ is the noise added by the diffusion process. To minimize the effect of unknown noise $\epsilon_{\text{ng}}$, it is desirable for $\eta_{1}+\eta_{2}+1$ to be sufficiently large. However, smaller values of $\eta_{1}$ and $\eta_{2}$ result in better consistency with the measurement $y$, as illustrated in Fig. 2. To balance this trade-off, we propose using a small $\eta_{1}+\eta_{2}$ during the early stages of sampling to fully exploit measurement information. As sampling progresses, $\eta_{1}+\eta_{2}$ should be gradually increased to suppress the impact of $\epsilon_{\text{ng}}$. The underlying intuition is that, in the early sampling stage, $x_{t}$ is still highly noisy and $a_{t}\approx 0$, so the influence of $\epsilon_{\text{ng}}$ is negligible even when $\eta_{1}+\eta_{2}$ is small. As shown in Fig. 3, $\tilde{x}^{\text{pm}}$ is initially more aligned with the degraded observation, but progressively shifts toward $\hat{x}_{0}^{\theta}$ as sampling evolves.
For a general degradation operator $H$, we recommend setting $\eta_{2}=ka_{t}/c_{t}$, where $k$ is a constant determined by the characteristics of the introduced noise. The rationale behind this design choice is further detailed in App. B.
Addressing non-differentiable measurements. For solving inverse problem with non-differentiable measurements such as JPEG restoration and quantization, the degraded images can be viewed as "noisy images" with unknown noise, modeled by $y=x+\epsilon_{y}$. In these scenarios, our proposed strategy naturally extends by treating the unknown degradations as implicit noise.
5 Experiments
5.1 Experimental setup
Dataset. We evaluate the effectiveness of our proposed approach on FFHQ $256\times 256$ (karras2019style, ) and ImageNet $256\times 256$ (deng2009imagenet, ). Following DAPS (zhang2024improving, ), we test on the same subset of 100 images for both datasets.
Pretrained models and baselines. We utilize the pre-trained checkpoint (chung2022diffusion, ) on the FFHQ dataset and the pre-trained checkpoint (dhariwal2021diffusion, ) on the Imagenet dataset. We compare our methods with the following baselines: DDNM (wang2022zero, ), DDRM (kawar2022denoising, ), $\Pi$GDM (song2023pseudoinverse, ), RedDiff (mardani2023variational, ), DAPS (zhang2024improving, ).
Metrics. Following previous work (chung2022diffusion, ; kawar2022denoising, ), we report Fréchet Inception Distance (FID) (heusel2017gans, ), Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable, ), Peak Signal-to-Noise Ratio (PSNR), and Structural SIMilarity index (SSIM).
Tasks. (1) We evaluate performance on the following linear inverse problems: super-resolution (pooling filter), deblurring (uniform kernel of size $25$), inpainting (box mask), inpainting (freeform mask), and colorization. (2) We consider two unknown noise types: salt-and-pepper noise ($10\%$ pixels set randomly to ±$\pm 1$) and periodic noise (sinusoidal pattern with amplitude $0.2$ and frequency $5$). (3) We address JPEG restoration with quality factors $\mathrm{QF}=2$ and $\mathrm{QF}=5$. (4). For quantization, we consider the challenging case of 2-bit quantization.
5.2 Ablation study
Figure 4: Ablation study of $\eta_{1}$ and $\eta_{2}$ on solving super-resolution and deblurring.
Ablation Study on $\eta_{1}$ and $\eta_{2}$.
We conduct ablation studies on parameters $\eta_{1}$ and $\eta_{2}$ using two inverse problems: super-resolution (noise-free, $\epsilon_{y}=0$) and deblurring (noisy, $\epsilon_{y}\sim\mathcal{N}(0,\sigma_{y}^{2}\mathbb{I})$). Results presented in Fig. 4 demonstrate that for noise-free super-resolution, the highest PSNR and lowest FID scores are achieved by setting $\eta_{2}=0$ and a negative $\eta_{1}=-0.4$. This indicates that appropriate "overshooting" enhances restoration quality. For the noisy deblurring task, negative $\eta_{2}$ similarly improves PSNR and lowers the FID score, further confirming the benefit of overshooting. Nevertheless, optimal performance is observed at $\eta_{2}=0.01$, suggesting that suitable regularization via $\eta_{2}$ significantly enhances reconstruction results. Additionally, we observe a performance drop as $\eta_{2}=0$ and $\eta_{1}\rightarrow 0$ (where $\tilde{x}_{0}^{\mathrm{bm}}\approx\tilde{x}_{0}^{\mathrm{DDNM}}$).
Table 2: Quantitative evaluation of solving noisy linear inverse problems with Gaussian noise (known variance, $\sigma_{y}=0.2$).
Task Method NFE FFHQ ImageNet
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
SR 8×8$\times$8 × DDNM 20 29.87 0.258 0.837 108.24 27.71 0.385 0.711 189.57
DDRM 20 29.53 0.269 0.834 114.73 27.40 0.397 0.704 192.43
$\Pi$GDM 100 21.21 0.632 0.384 369.61 19.66 0.604 0.284 303.60
RED-Diff 100 24.06 0.604 0.604 320.15 23.22 0.61 0.522 310.29
DAPS 100 18.71 0.622 0.370 328.65 17.18 0.658 0.275 309.68
MAS (ours) 20 29.88 0.254 0.838 107.67 27.71 0.377 0.712 185.43
Deblurring DDNM 20 21.31 0.571 0.173 175.71 19.98 0.592 0.134 295.76
DDRM 20 29.25 0.271 0.823 121.00 27.54 0.405 0.704 213.13
$\Pi$GDM 100 5.15 0.846 0.004 493.09 - - - -
RED-Diff 100 5.20 0.845 0.004 502.66 5.20 0.804 0.004 457.13
DAPS 100 19.01 0.616 0.365 244.10 16.60 0.668 0.231 307.79
MAS (ours) 20 29.62 0.267 0.827 119.43 27.88 0.408 0.710 218.51
Inpaint (Freeform) DDNM 20 25.99 0.191 0.887 117.12 27.04 0.218 0.854 98.60
DDRM 20 26.53 0.181 0.903 93.73 27.28 0.205 0.855 75.71
$\Pi$GDM 100 24.88 0.396 0.504 129.45 25.01 0.352 0.513 119.64
RED-Diff 100 25.22 0.404 0.517 174.56 25.47 0.373 0.538 150.64
DAPS 100 18.99 0.471 0.330 148.37 19.43 0.419 0.385 138.87
MAS (ours) 20 27.20 0.157 0.913 92.56 27.92 0.177 0.871 77.61
Inpaint (Box) DDNM 20 27.87 0.165 0.880 94.27 24.88 0.247 0.826 168.99
DDRM 20 28.36 0.159 0.896 84.76 24.87 0.236 0.827 151.55
$\Pi$GDM 100 25.72 0.375 0.509 153.81 23.44 0.377 0.486 151.99
RED-Diff 100 25.87 0.382 0.522 186.75 24.19 0.409 0.518 239.49
DAPS 100 19.86 0.437 0.338 163.03 18.03 0.448 0.354 212.95
MAS (ours) 20 28.77 0.135 0.904 83.79 24.75 0.213 0.838 143.01
Colorization DDNM 20 28.41 0.216 0.937 113.04 27.08 0.315 0.877 162.19
DDRM 20 28.34 0.220 0.930 108.50 27.19 0.285 0.876 124.94
$\Pi$GDM 100 23.51 0.552 0.409 185.22 22.57 0.513 0.453 216.30
RED-Diff 100 24.50 0.557 0.403 206.85 23.67 0.537 0.462 274.81
DAPS 100 21.71 0.491 0.479 145.99 20.45 0.492 0.531 185.29
MAS (ours) 20 28.43 0.205 0.941 112.94 27.39 0.282 0.893 147.07
5.3 Image restoration
Inverse problems with Gaussian noise (known variance). Quantitative results for inverse problems with Gaussian noise of known variance are shown in Table 2. MAS consistently demonstrates superior performance across most tasks, notably achieving significantly lower FID scores.
Table 3: Quantitative evaluation of solving linear inverse problems with non-Gaussian noise (unknown strength).
Task Method Salt peper noise Periodic noise
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
SR 8×8$\times$8 × DDNM 20.23 0.592 0.597 368.76 24.07 0.448 0.725 265.62
DDRM 20.26 0.592 0.602 360.79 24.11 0.451 0.729 268.48
$\Pi$GDM 20.31 0.575 0.616 338.50 24.04 0.432 0.726 222.93
RED-Diff 20.72 0.590 0.656 400.16 24.10 0.478 0.697 315.58
MAS (ours) 25.68 0.352 0.783 124.66 25.63 0.364 0.768 132.11
Deblurring DDNM 5.11 0.845 0.004 500.61 6.30 0.764 0.049 442.27
DDRM 5.11 0.845 0.004 499.67 6.30 0.764 0.049 446.33
$\Pi$GDM 5.10 0.845 0.004 497.12 6.30 0.769 0.048 451.63
RED-Diff 5.11 0.844 0.004 501.76 7.17 0.755 0.073 506.21
MAS (ours) 27.10 0.314 0.788 111.51 23.47 0.409 0.742 138.48
Inverse Problems with Non-Gaussian Noise (Unknown Strength). Quantitative evaluations for linear inverse problems with unknown, non-Gaussian noise are presented in Table 3. MAS outperforms baseline methods across all metrics, highlighting the effectiveness of our approach in handling unknown noise conditions.
Figure 5: Results on JPEG (QF=5) and quantization restoration.
Inverse problems with non-differentiable measurements. MAS is also capable of solving inverse problem with non-differentiable measurements, such as JPEG restoration and quantization. Results in Table 4 and Fig. 5 show that MAS achieves state-of-the-art performance without relying on the forward operator or knowledge of the degradation strength.
Table 4: Quantitative evaluation of solving JPEG restoration and Quantization. We set $k=1.0$ for $\mathrm{QF}=5$ and $k=3.0$ for $\mathrm{QF}=2$, and $k=0.5$ for quantization. Both $\Pi$GDM and MAS use $\mathrm{NFE}=20$, which yields the best performance (among $\mathrm{NFE}=20$ and $\mathrm{NFE}=100$). Notably, our method (MAS) does not require access to the forward operator or the strength of degration.
Method JPEG Restoration (QF = 5) JPEG Restoration (QF = 2) Quantization (number of bits = 2)
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
$\Pi$GDM 32.25 0.203 0.887 130.57 29.35 0.270 0.838 147.93 35.15 0.150 0.922 112.40
MAS (ours) 32.31 0.225 0.898 101.26 29.73 0.277 0.862 114.83 34.98 0.149 0.937 69.61
6 Related work
Diffusion models have also been successfully applied to linear inverse problems, including, compressed-sensing MRI (CS-MRI), and computed tomography (CT) (kadkhodaie2021stochastic, ; song2020score, ; chung2022improving, ; kawar2022denoising, ; song2021solving, ). They have also been extended to non-linear inverse problems such as Fourier phase retrieval, nonlinear deblurring, HDR, and JPEG restoration (chung2022diffusion, ; song2023pseudoinverse, ; chung2023parallel, ; mardani2023variational, ). RED-diff (mardani2023variational, ) leverages stochastic optimization to approximate the true posterior. Besides, Chung et al. (chung2023decomposed, ) combine the diffusion sampling and Krylov subspace methods, which negates the need to compute the manifold-constrained gradient. InverseBench (zheng2025inversebench, ) presents a benchmark for critical scientific applications, which present structural challenges that differ significantly from natural image restoration tasks.
In recent years, diffusion bridge models have been successfully applied to various inverse problems (liu2023i2sb, ; zhou2023denoising, ; albergo2023stochastic2, ). These models are more task-specific and directly build probabilistic paths from corrupted to clean images, showing strong performance. Exploring alignment techniques within diffusion bridges is a promising direction for further improvement.
7 Conclusion
MAS improves upon existing methods by explicitly aligning the sampling process with measurement data, offering a broader optimization perspective that generalizes approaches like DDNM and DAPS. Beyond the noise-free case, MAS can be extended to: (1) known Gaussian noise, (2) unknown or non-Gaussian noise through adaptive parameterization, and (3) non-differentiable degradations (e.g., JPEG) by decoupling the forward operator from sampling. Extensive experiments show that MAS consistently outperforms state-of-the-art methods across a wide range of inverse problems.
Limitations. (i) While MAS can handle non-differentiable measurements like JPEG restoration, it does not support general non-linear inverse problems. (ii) Tweedie’s formula boys2023tweedie could be used to improve the MAS prior by better estimating the covariance of $p(x_{0}\mid x_{t})$. Also, enhancing the likelihood model $p(y\mid\tilde{x}_{0})$ may further improve performance.
Appendix A Proofs
A.1 Proof of Prop. 3.1.
Proof.
Write
$\Sigma_{y}\;\coloneqq\;r_{t}^{2}\,HH^{\mathsf{T}}+\sigma_{y}^{2}\,\mathbb{I}_{m},\qquad\Sigma_{0}\;\coloneqq\;\rho_{t}^{2}\,\mathbb{I}_{d},\qquad\mu_{0}\;\coloneqq\;\hat{x}_{0}^{\theta}.$ (21)
Because both the likelihood $p(y\mid\tilde{x}_{0})=\mathcal{N}\!\bigl{(}H\tilde{x}_{0},\Sigma_{y}\bigr{)}$ and the prior $p(\tilde{x}_{0}\mid\hat{x}_{0}^{\theta})=\mathcal{N}\!\bigl{(}\mu_{0},\Sigma_{0}\bigr{)}$ are multivariate Gaussians, their product is proportional to another Gaussian. Using Bayes’ theorem and taking logarithms,
$\log p(\tilde{x}_{0}\mid y,\hat{x}_{0}^{\theta})=-\tfrac{1}{2}\!\Bigl{[}(y-H\tilde{x}_{0})^{\mathsf{T}}\Sigma_{y}^{-1}(y-H\tilde{x}_{0})+(\tilde{x}_{0}-\mu_{0})^{\mathsf{T}}\Sigma_{0}^{-1}(\tilde{x}_{0}-\mu_{0})\Bigr{]}+C.$ (22)
Expand the quadratic terms and collect everything that depends on $\tilde{x}_{0}$:
$\log p(\tilde{x}_{0}\mid y,\hat{x}_{0}^{\theta})=-\tfrac{1}{2}\tilde{x}_{0}^{\mathsf{T}}\bigl{(}\Sigma_{0}^{-1}+H^{\mathsf{T}}\Sigma_{y}^{-1}H\bigr{)}\tilde{x}_{0}+\tilde{x}_{0}^{\mathsf{T}}\bigl{(}\Sigma_{0}^{-1}\mu_{0}+H^{\mathsf{T}}\Sigma_{y}^{-1}y\bigr{)}+C^{\prime}.$ (23)
The exponent is a (shifted) quadratic form in $\tilde{x}_{0}$; completing the square yields
$\log p(\tilde{x}_{0}\mid y,\hat{x}_{0}^{\theta})=-\tfrac{1}{2}\!\bigl{(}\tilde{x}_{0}-\mu_{\mathrm{post}}\bigr{)}^{\mathsf{T}}\Sigma_{\mathrm{post}}^{-1}\bigl{(}\tilde{x}_{0}-\mu_{\mathrm{post}}\bigr{)}+C^{\prime\prime},$ (24)
where
$\boxed{\;\Sigma_{\mathrm{post}}=\bigl{(}\Sigma_{0}^{-1}+H^{\mathsf{T}}\Sigma_{y}^{-1}H\bigr{)}^{-1}\;}\quad\text{and}\quad\boxed{\;\mu_{\mathrm{post}}=\Sigma_{\mathrm{post}}\,\bigl{(}\Sigma_{0}^{-1}\mu_{0}+H^{\mathsf{T}}\Sigma_{y}^{-1}y\bigr{)}\;}.$ (25)
Substituting $\Sigma_{0}^{-1}=\rho_{t}^{-2}\mathbb{I}_{d}$ and $\Sigma_{y}^{-1}=(r_{t}^{2}HH^{\mathsf{T}}+\sigma_{y}^{2}\mathbb{I}_{m})^{-1}$ gives precisely the expressions in (9). Hence the posterior distribution is the multivariate Gaussian stated in the proposition. ∎
A.2 Proof of effieient linear solves in Eq. 13
Lemma A.1.
Let $H\in\mathbb{R}^{m\times d}$ have (thin) singular–value decomposition $H=U\Sigma V^{\mathsf{T}}$ with orthogonal $U\in\mathbb{R}^{m\times m}$, $V\in\mathbb{R}^{d\times d}$ and $\Sigma=\operatorname{diag}(s_{1},\dots,s_{r})\in\mathbb{R}^{m\times d}$, where $r=\operatorname{rank}(H)$ and $s_{1}\geq\cdots\geq s_{r}>0$. For any scalars $\eta_{1}\geq 0$ and $\eta_{2}>0$ define
$W\coloneqq\eta_{1}HH^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m},\qquad Y\coloneqq\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H.$
Then
[…]
(When $i>r$ we set $s_{i}\!=\!0$.)
Proof.
(i) Inverting $W$. Using the SVD,
$W=\eta_{1}U\Sigma\Sigma^{\mathsf{T}}U^{\mathsf{T}}+\eta_{2}U\mathbb{I}_{m}U^{\mathsf{T}}=U\!\bigl{(}\eta_{1}\Sigma\Sigma^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m}\bigr{)}U^{\mathsf{T}}.$
Because $U$ is orthogonal, $W^{-1}$ is obtained by inverting the diagonal middle matrix: $(\eta_{1}\Sigma\Sigma^{\mathsf{T}}+\eta_{2}\mathbb{I}_{m})^{-1}=\operatorname{diag}\!\bigl{(}\frac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr{)}_{i=1}^{m}.$ Substituting yields the first identity in (26).
(ii) Inverting $Y$. Write
$Y=\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H=V\Sigma^{\mathsf{T}}U^{\mathsf{T}}\bigl{[}U\operatorname{diag}\!\bigl{(}\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr{)}U^{\mathsf{T}}\bigr{]}U\Sigma V^{\mathsf{T}}+\mathbb{I}_{d},$
and simplify with $U^{\mathsf{T}}U=\mathbb{I}_{m}$:
$Y=V\!\Bigl{[}\Sigma^{\mathsf{T}}\!\operatorname{diag}\!\bigl{(}\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr{)}\!\Sigma+\mathbb{I}_{d}\Bigr{]}V^{\mathsf{T}}.$
Because $\Sigma^{\mathsf{T}}\!\operatorname{diag}\!\bigl{(}\tfrac{1}{\eta_{1}s_{i}^{2}+\eta_{2}}\bigr{)}\Sigma$ is diagonal with $i^{\text{th}}$ entry $\tfrac{s_{i}^{2}}{\eta_{1}s_{i}^{2}+\eta_{2}}$, the bracketed matrix is diagonal and hence trivial to invert, giving the second identity in (26). ∎
A.3 Proof of Eq. 14
Proposition A.2.
Let $H\in\mathbb{R}^{m\times d}$, $\eta_{1}\geq 0$ and $\eta_{2}>0$. Define
$W\;=\;\eta_{1}\,HH^{\mathsf{T}}+\eta_{2}\,\mathbb{I}_{m},\qquad Y\;=\;\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H.$
For any $y\in\mathbb{R}^{m}$ and $\hat{x}_{0}^{\theta}\in\mathbb{R}^{d}$ consider the strictly convex quadratic
$\mathcal{L}(\tilde{x}_{0})\;=\;\bigl{\|}\tilde{x}_{0}-\hat{x}_{0}^{\theta}\bigr{\|}_{2}^{2}\;+\;\bigl{\|}y-H\tilde{x}_{0}\bigr{\|}_{W^{-1}}^{2},\qquad\|v\|_{W^{-1}}^{2}=v^{\mathsf{T}}W^{-1}v.$
Its unique minimiser is
$\tilde{x}_{0}^{\mathrm{pm}}\;=\;Y^{-1}\!\bigl{[}\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y\bigr{]}.$ (27)
Proof.
Expand $\mathcal{L}$ and take its gradient:
[…]
Setting $\nabla_{\tilde{x}_{0}}\mathcal{L}=0$ gives the normal equation
$\bigl{(}\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H\bigr{)}\tilde{x}_{0}\;=\;\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y,\qquad\text{that is, }Y\,\tilde{x}_{0}=\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y.$
Because $\eta_{2}>0$ implies $W\succ 0$, we have $W^{-1}\succ 0$ and hence $Y=\mathbb{I}_{d}+H^{\mathsf{T}}W^{-1}H\succ 0$; thus $Y$ is invertible and (27) follows.
Finally, the Hessian of $\mathcal{L}$ is $2Y\succ 0$, so $\mathcal{L}$ is strictly convex and the stationary point (27) is indeed its unique global minimiser. ∎
A.4 Proof of Remark 2.
Proof.
We start with the posterior expectation
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\left(\frac{1}{\rho_{t}^{2}}\mathbb{I}+\frac{1}{r_{t}^{2}}H^{\dagger}H\right)^{-1}\left(\frac{1}{\rho_{t}^{2}}\hat{x}_{0}^{\theta}+\frac{1}{r_{t}^{2}}H^{\dagger}y\right).$ (28)
We want to examine the limit as
$\frac{r_{t}}{\rho_{t}}\rightarrow 0,$ (29)
which implies that $r_{t}^{2}/\rho_{t}^{2}\rightarrow 0$. To make this explicit, set
$\epsilon=\frac{r_{t}^{2}}{\rho_{t}^{2}},$ (30)
so that $\epsilon\to 0$. Then we can write
$\frac{1}{\rho_{t}^{2}}=\frac{1}{r_{t}^{2}}\,\epsilon,$ (31)
and the expression becomes
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\left(\frac{\epsilon}{r_{t}^{2}}\,\mathbb{I}+\frac{1}{r_{t}^{2}}H^{\dagger}H\right)^{-1}\left(\frac{\epsilon}{r_{t}^{2}}\hat{x}_{0}^{\theta}+\frac{1}{r_{t}^{2}}H^{\dagger}y\right).$ (32)
Factor out $1/r_{t}^{2}$ from both terms:
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\left(\frac{1}{r_{t}^{2}}\Bigl{(}\epsilon\,\mathbb{I}+H^{\dagger}H\Bigr{)}\right)^{-1}\left(\frac{1}{r_{t}^{2}}\Bigl{(}\epsilon\,\hat{x}_{0}^{\theta}+H^{\dagger}y\Bigr{)}\right).$ (33)
The $1/r_{t}^{2}$ factors cancel (since the inverse of $aM$ is $M^{-1}/a$), yielding
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\left(\epsilon\,\mathbb{I}+H^{\dagger}H\right)^{-1}\left(\epsilon\,\hat{x}_{0}^{\theta}+H^{\dagger}y\right).$ (34)
To analyze the limit as $\epsilon\to 0$, decompose the space into two orthogonal components:
  • The range (or row space) of $H$, on which $H^{\dagger}H$ acts as the identity.
  • Its nullspace, on which $H^{\dagger}H$ is zero.
Let
$P=H^{\dagger}H,$ (35)
which is the orthogonal projection onto the row space of $H$. Then any vector $v$ can be decomposed as
$v=Pv+(I-P)v.$ (36)
Notice that $H^{\dagger}y$ lies in the row space (i.e. $P\,H^{\dagger}y=H^{\dagger}y$) and that $\hat{x}_{0}^{\theta}$ can be decomposed as
$\hat{x}_{0}^{\theta}=P\hat{x}_{0}^{\theta}+(I-P)\hat{x}_{0}^{\theta}.$ (37)
Since the eigenvalues of $P$ are 0 and 1, the matrix $\epsilon\,I+P$ has eigenvalues $\epsilon$ (on the nullspace of $P$) and $1+\epsilon$ (on the row space). Hence, its inverse acts as:
  • Multiplication by $1/\epsilon$ on the nullspace,
  • Multiplication by $1/(1+\epsilon)$ on the row space.
Thus, we have
$\left(\epsilon\,I+P\right)^{-1}\left(H^{\dagger}y+\epsilon\,\hat{x}_{0}^{\theta}\right)=\frac{1}{1+\epsilon}\bigl{(}H^{\dagger}y+\epsilon\,P\hat{x}_{0}^{\theta}\bigr{)}+\frac{1}{\epsilon}\bigl{(}\epsilon\,(I-P)\hat{x}_{0}^{\theta}\bigr{)}.$ (38)
Simplify this to obtain
$\frac{1}{1+\epsilon}\,H^{\dagger}y+\frac{\epsilon}{1+\epsilon}\,P\hat{x}_{0}^{\theta}+(I-P)\hat{x}_{0}^{\theta}.$ (39)
Now, taking the limit as $\epsilon\to 0$:
  • $\frac{1}{1+\epsilon}\to 1$,
  • $\frac{\epsilon}{1+\epsilon}\to 0$.
Therefore, the limit becomes
$\lim_{\epsilon\to 0}\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=H^{\dagger}y+(I-P)\hat{x}_{0}^{\theta}.$ (40)
Recalling that $P=H^{\dagger}H$, we rewrite this as
$H^{\dagger}y+\hat{x}_{0}^{\theta}-H^{\dagger}H\,\hat{x}_{0}^{\theta}=\hat{x}_{0}^{\theta}+H^{\dagger}\bigl{(}y-H\hat{x}_{0}^{\theta}\bigr{)}.$ (41)
Thus, in the limit where $\frac{r_{t}}{\rho_{t}}\to 0$, we indeed have
$\mathbb{E}[\tilde{x}_{0}\mid\hat{x}_{0}^{\theta},y]=\hat{x}_{0}^{\theta}+H^{\dagger}\bigl{(}y-H\hat{x}_{0}^{\theta}\bigr{)}.$ (42)
This shows that, as the relative measurement noise $r_{t}$ becomes much smaller compared to the prior uncertainty $\rho_{t}$, the posterior expectation is the projection of $\hat{x}_{0}^{\theta}$ onto the subspace $\{x:Hx=y\}$. ∎
Appendix B Additional method details
B.1 Addressing Gaussian noise
Consider noisy image restoration problems in the form of $y=Hx+\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\epsilon_{y}}$, where $\epsilon_{y}$ is the added noise. Then the measurement $y$ can be decomposed to the sum of clean measurement $y^{\text{clean}}\coloneqq Hx$ and measurement noise $\epsilon_{y}$. Calculating posterior mean $\tilde{x}_{0}^{\text{pm}}$ leads to:
$\tilde{x}_{0}^{\text{pm}}$ $=Y^{-1}[\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (43)
$=\hat{x}_{0}^{\theta}+(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+Y^{-1}H^{T}W^{-1}y$ (44)
where ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{T}W^{-1}\epsilon_{y}}$ is the extra noise introduced into $\tilde{x}_{0}^{\text{pm}}$ and will be further introduced into $x_{t-\Delta t}$.
To address Gaussian noise with known variance, we modify Eq. 12 and Eq. 3 as:
$\tilde{x}_{0}^{\text{pm}}=\hat{x}_{0}^{\theta}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}[(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (45)
$x_{t-\Delta t}\sim\mathcal{N}(a_{t}\tilde{x}_{0}^{\text{pe}}(t,x,y)+b_{t}x_{t},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Phi}_{t}\mathbb{I})$ (46)
Then $\tilde{x}_{0}^{\text{pm}}$ is:
$\tilde{x}_{0}^{\text{pm}}$ $=\hat{x}_{0}^{\theta}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}[(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (47)
$=\underbrace{\hat{x}_{0}^{\theta}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}}(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+Y^{-1}H^{T}W^{-1}y^{\text{clean}}}_{\coloneqq\;\tilde{x}_{0}^{\text{clean}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma_{t}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ (49)
Then the iteration of the sampling process is:
$x_{t-\Delta t}$ $=a_{t}\tilde{x}_{0}^{\text{pm}}(t,x,y)+b_{t}x_{t}+\epsilon_{\text{new}},\quad\epsilon_{\text{new}}\sim\mathcal{N}(0,\Phi_{t})$ (50)
$=a_{t}\tilde{x}_{0}^{\text{clean}}+b_{t}x_{t}+\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{t}\sigma_{y}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}}_{\coloneqq\epsilon_{\text{intro}}}+\epsilon_{\text{new}}$ (51)
Suppose $\Sigma_{t}=V\mathrm{diag}\{\lambda_{t1},\cdots,\lambda_{td}\}V^{T}$ $\Phi_{t}=Vdiag\{\gamma_{t1},\cdots\gamma_{td}\}V^{T}$. Then the introduced noise $\epsilon_{\text{intro}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{t}\sigma_{y}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ is still a Gaussian distribution: $\epsilon_{\text{intro}}\sim\mathcal{N}(0,VD_{t}V^{T})$, with $D_{t}=diag\{d_{t1},\cdots,d_{td}\}$:
$d_{ti}\;=\;\begin{cases}\displaystyle\frac{a_{t}^{2}\,\sigma_{y}^{2}\,s_{i}^{2}\lambda_{ti}^{2}}{\bigl{[}(\eta_{1}+1)\,s_{i}^{2}+\eta_{2}\bigr{]}^{2}},&s_{i}\neq 0,\\[8.0pt] 0,&s_{i}=0,\end{cases}$ (52)
The choice of and $\Phi_{t}$ need to ensure the total noise injected to $x_{t-\Delta t}$ conforms the iteration in Eq. 3.
$\epsilon_{\text{new}}+\epsilon_{\text{intro}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$ (53)
To construct $\epsilon_{\text{new}}$, we define a new diagonal matrix $\Gamma_{t}(=diag\{\gamma_{t1},\cdots\gamma_{td}\})$:
$\gamma_{ti}\;=\;\begin{cases}\displaystyle c_{t}^{2}-\frac{a_{t}^{2}\,\sigma_{y}^{2}\,s_{i}^{2}\lambda_{ti}^{2}}{\bigl{[}(\eta_{1}+1)\,s_{i}^{2}+\eta_{2}\bigr{]}^{2}},&s_{i}\neq 0,\\[8.0pt] c_{t}^{2},&s_{i}=0,\end{cases}$ (54)
Now we can yield $\epsilon_{\text{new}}$ by sampling from $\mathcal{N}(0,V\Gamma_{t}V^{T})$ to ensure that $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. We need to make sure $\lambda_{ti}$ guaratees the noise level of the introduced noise do not exceed the pre-defined noise level $c_{t}$, we also hope $\lambda_{t_{i}}$ as close as $1$ as possible. Therefore,
$\lambda_{ti}\;=\;\begin{cases}\displaystyle 1,&c_{t}\geq\frac{a_{t}\,\sigma_{y}\,s_{i}}{(\eta_{1}+1)\,s_{i}^{2}+\eta_{2}},\\ \frac{c_{t}((\eta+1)s_{i}^{2}+\eta_{2})}{a_{t}\sigma_{y}s_{i}},&c_{t}<\frac{a_{t}\,\sigma_{y}\,s_{i}}{(\eta_{1}+1)\,s_{i}^{2}+\eta_{2}},\\[8.0pt] 1,&s_{i}=0.\end{cases}$ (55)
B.2 Addressing unknown noise and non-differentiable measurements
As the measurement noise $\epsilon_{y}$ is non-Gaussian or unknown, it’s difficult to ensure the total noise introduced in $x_{t-\Delta t}$ is $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. In this case, we calculate $\tilde{x}_{0}^{\text{pm}}$ using Eq. 12 and update $x_{t-\Delta t}$ using Eq. 3. Then $\tilde{x}_{0}^{\text{pm}}$ is:
$\tilde{x}_{0}^{\text{pm}}$ $=\hat{x}_{0}^{\theta}+[(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+H^{\mathsf{T}}W^{-1}y]$ (56)
$=\underbrace{\hat{x}_{0}^{\theta}+(Y^{-1}-\mathbb{I}_{d})\hat{x}_{0}^{\theta}+Y^{-1}H^{T}W^{-1}y^{\text{clean}}}_{\coloneqq\;\tilde{x}_{0}^{\text{clean}}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}$ (58)
where ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Y^{-1}H^{T}W^{-1}\epsilon_{y}}$ is the extra noise introduced into $\tilde{x}_{0}^{\text{pm}}$ and will be further introduced into $x_{t-\Delta t}$:
$x_{t-\Delta t}$ $=a_{t}\tilde{x}_{0}^{\text{pm}}+b_{t}x_{t}+\epsilon_{\text{new}},$ (59)
$=a_{t}\tilde{x}_{0}^{\text{clean}}+b_{t}x_{t}+\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}a_{t}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}}}_{\coloneqq\;\epsilon_{\text{intro}}}+\epsilon_{\text{new}},$
where $\epsilon_{\text{new}}$ the noise added by diffusion process, which should be specifically designed to ensure $x_{t-\Delta t}$ is sampled from correct distribution as in Eq. 3, i.e., the total noise $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. However, as $\epsilon_{y}$ is unknown noise, we have no information about the introduced noise $\epsilon_{\text{intro}}$. To solve this problem, we made the following principles: (i) despite that fact that we cannot guarantee $\epsilon_{\text{intro}}+\epsilon_{\text{new}}\sim\mathcal{N}(0,c_{t}^{2}\mathbb{I})$, we still hope $\epsilon_{\text{intro}}+\epsilon_{\text{new}}$ is as close to $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$ as possible; (ii) small $\eta_{1}$ and $\eta_{2}$ are helpful to maximize the alignment to measurement $y$. Notably,
$\epsilon_{\text{intro}}$ $=a_{t}Y^{-1}H^{\mathsf{T}}W^{-1}\epsilon_{y}$ (60)
$=a_{t}V\mathrm{diag}\left(\frac{s_{i}}{(\eta_{1}+1)s_{i}^{2}+\eta_{2}}\right)U^{T}\epsilon_{y}$ (61)
$\eta_{1}$ and $\eta_{2}$ are two variables that control the noise level of $\epsilon_{\text{intro}}$. In the implementation, we still sample $\epsilon_{\text{new}}$ from Gaussian distribution $\mathcal{N}(0,c_{t}^{2}\mathbb{I})$. Then the problem becomes how to select $\eta_{1}$ and $\eta_{2}$ to meet the above 2 principles. For common image restoration tasks like SR, Deblurring, inpainting, Colorization, The maximum eigenvalue value $s_{\max}=\max\{s_{i}\}<=1$. Therefore, adjusting $\eta_{2}$ is more likely to reduce the strength of $\epsilon_{\text{intro}}$.
Fix a tiny baseline $\eta_{1}\in[-0.4,0)\cup(0,0.1)$, for example, we can choose $\eta_{2}(t)=ka_{t}/c_{t}$, where $k$ is a constant wihch depends on the measurement noise $\epsilon_{y}$. This meets Principle (i) by bounding the extra noise, and Principle (ii) by keeping both $\eta_{1}$, $\eta_{2}$ small enough to preserve measurement alignment.
Appendix C Limitations
There are some limitations that deserve further study.
  • Theoretical extension. While MAS can, in principle, be generalized to nonlinear inverse problems, explicitly formulating the likelihood term $p(y\mid\tilde{x}_{0})$ becomes challenging. Developing effective sampling techniques under this setting is a promising direction for future research.
  • Improved prior modeling. Boys et al. [2] leverage Tweedie’s formula to compute the covariance of the Gaussian projection of $p(x_{0}\mid x_{t})$, which could be incorporated to enhance the MAS prior in Eq. 7a. Moreover, improved modeling of the likelihood $p(y\mid\tilde{x}_{0})$ may further boost the performance of MAS.
Appendix D Impact Statement
Our method can improve image restoration under challenging noise and degradation conditions, which may benefit applications in medical imaging, scientific visualization, cultural heritage preservation, and general photography. However, it is important to note that as with many generative and restoration models, our method could be misused for malicious image manipulation.
Appendix E Pytorch-like code implementation
Here we provide a basic PyTorch-Like implementation of the calculation of $\tilde{x}_{0}^{\text{pm}}$ in Eq. 12, shown in LABEL:lst:mas_code.
1@torch.no_grad()
2def mas(
3 H, x0_hat, y,
4 eta_1=-0.2, eta_2=0.0
5):
6 bs, _, H_img, W_img = x0_hat.shape
7 x0_hat = x0_hat.view(bs, -1)
8 y = y.view( bs, -1) # measurement dim m
9 ut_y = H.Ut(y) # (bs, m)
10 singulars = H.singulars() # (m,)
11 nz = singulars > 0 # boolean mask
12 scale1 = 1.0 / (singulars[nz] ** 2 * eta_1 + eta_2)
13 ut_y[:, nz] = ut_y[:, nz] * scale1 # broadcasting OK
14 u_y = H.U(ut_y) # (bs, m)
15 rhs = x0_hat + H.Ht(u_y) # (bs, d)
16
17 vt_rhs = H.Vt(rhs) # (bs, d)
18 scale2 = 1.0 / (1.0 + singulars[nz]**2 / (singulars[nz]**2 * eta_1 + eta_2))
19 vt_rhs[:, nz] = vt_rhs[:, nz] * scale2
20 x0_pm = H.V(vt_rhs) # (bs, d)
21 x0_pm = x0_pm.view(bs, 3, H_img, W_img)
22 return x0_pm
Listing 1: PyTorch-like implementation of the calculation of $\tilde{x}_{0}^{\text{pm}}$ in Eq. 12.
Appendix F Experimental details
F.1 Details of the degradation operators
Super-resolution. We use the averagepooling downsampler as the forward operator.
Deblurring. For deblurring experiments, We use uniform blur kernel to to implement blurring operation.
Figure 6: Restoration of inpaint (freeform) using MAS.
Inpaint (freeform). The freeform mask covers approximately $19.3\%$ of the total image area. Examples of restoration are shown in Fig. 6.
Inpaint (box). We use a fixed square mask of size $128\times 128$ pixels placed at the center of the image.
Colorization. We simulate grayscale degradation by applying a fixed linear transformation to each pixel using the matrix $[\frac{1}{3},\frac{1}{3},\frac{1}{3}]$, replacing each RGB pixel with its average intensity.
F.2 Details of the baseline models
Sampler. Most experiments on diffusion models leverage DDIM [21] sampling.
DDRM [14]. $\eta_{B}=1.0$, $\eta=0.85$ with DDIM sampler, as advised in the original paper.
$\Pi$GDM [22]. $\eta=1.0$, with DDIM sampler, as advised in the original paper.
Reddiff [18]. $\lambda=0.25$, with DDIM sampler, as advised in the original paper.
DDNM [25]. $\eta=0.85$, with DDIM sampler, as advised in the original paper.
DAPS [27]. $\tau=0.01$, 100 steps of MCMC, with EDM sampler, as advised in the original paper.
Appendix G Additional results
G.1 Computational time
The computational time of MAS on solving inverse problems is shown in Table 5. Our model achieves similar efficiency to DDNM and DDRM, demonstrating that MAS introduces minimal overhead while maintaining competitive runtime performance.
Table 5: Sampling time (Sec) per image of MAS on deblurring and super-resolution with FFHQ 256, evaluated using a single NVIDIA A6000 48G GPU. We set NFE=20 for all of the methods.
Method MAS $\Pi$GDM DDNM DDRM RED-Diff
Deblurring 0.128 0.278 0.127 0.127 0.119
SR (8×8$\times$8 ×) 0.131 0.282 0.131 0.131 0.125
G.2 Comparison between posterior sampling and posterior mean
We show the comparison between posterior sampling Eq. 11 and posterior mean Eq. 12 in Fig. 7. The experiments results demonstrate that posterior mean achieve the best results. posterior sampling will introduce additional noise to the final restoration.
Figure 7: The comparison between posterior sampling in Eq. 11 and posterior mean Eq. 12 on deblurring task. We set $\eta_{1}=-0.2$ and $\eta_{2}=0$.
Table 6: Quantitative evaluation of solving super-resolution and deblurring. We set $\eta_{1}=-0.2$ and $\eta_{2}=0$ for all of the tasks.
Strength DDNM MAS
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
4$\times$ 36.77 0.080 0.930 64.60 36.71 0.075 0.929 43.81
8$\times$ 32.28 0.180 0.863 90.86 32.46 0.170 0.866 74.60
16$\times$ 28.29 0.282 0.800 107.95 28.50 0.272 0.805 101.52
32$\times$ 24.79 0.374 0.750 131.64 25.01 0.363 0.755 130.78
64$\times$ 21.69 0.449 0.720 152.05 21.86 0.441 0.724 151.10
128$\times$ 19.10 0.497 0.705 145.61 19.19 0.496 0.707 147.10
G.3 Noise-free inverse problem
Quantitative comparisons for solving noise-free inverse problems are presented in Table 7. The results demonstrate that our model consistently outperforms DDNM, achieving significantly lower FID scores.
We include additional quantitative evaluation of solving super-resolution and deblurring under different strength of degradation, see Table 8 and Table 9 for details. MAS constantly outperforms DDNM under different strength of degradation.
Table 7: Quantitative evaluation of solving noise-free linear inverse problems. We set $\eta_{1}=-0.2$, $\eta_{2}=0$ for all of the tasks.
Task Method NFE FFHQ ImageNet
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
SR 8×8$\times$8 × DDNM 20 32.28 0.180 0.863 90.76 28.13 0.294 0.693 150.42
DDRM 20 32.33 0.180 0.865 87.54 29.22 0.296 0.735 142.62
$\Pi$GDM 20 32.60 0.177 0.871 87.35 29.22 0.302 0.738 149.87
RED-Diff 20 31.26 0.286 0.823 160.28 28.64 0.362 0.714 254.99
MAS (ours) 20 32.46 0.170 0.866 76.80 29.11 0.282 0.732 135.17
DAPS 100 26.05 0.269 0.690 101.03 23.17 0.385 0.545 161.43
$\Pi$GDM 100 32.03 0.149 0.856 68.79 28.37 0.242 0.705 103.24
RED-Diff 100 33.52 0.191 0.887 113.99 29.87 0.311 0.752 157.61
Deblurring DDNM 20 41.60 0.063 0.967 44.82 37.37 0.107 0.912 51.50
DDRM 20 41.61 0.060 0.968 43.22 37.96 0.096 0.921 47.39
$\Pi$GDM 20 42.34 0.047 0.974 38.71 37.45 0.103 0.918 58.65
RED-Diff 20 23.82 0.446 0.665 185.93 23.13 0.501 0.560 314.05
MAS (ours) 20 42.34 0.041 0.971 33.27 38.25 0.085 0.918 37.09
DAPS 100 25.28 0.300 0.677 110.67 22.78 0.438 0.519 200.86
$\Pi$GDM 100 44.92 0.024 0.983 23.40 40.63 0.049 0.947 20.05
RED-Diff 100 28.51 0.339 0.814 167.57 26.47 0.507 0.676 302.68
Table 8: Quantitative evaluation of solving super-resolution of different strength of degradation. We set $\eta_{1}=-0.4$ and $\eta_{2}=0$ for all of the tasks.
Strength DDNM MAS
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
4$\times$ 36.77 0.080 0.930 64.60 36.71 0.075 0.929 43.81
8$\times$ 32.28 0.180 0.863 90.86 32.46 0.170 0.866 74.60
16$\times$ 28.29 0.282 0.800 107.95 28.50 0.272 0.805 101.52
32$\times$ 24.79 0.374 0.750 131.64 25.01 0.363 0.755 130.78
64$\times$ 21.69 0.449 0.720 152.05 21.86 0.441 0.724 151.10
128$\times$ 19.10 0.497 0.705 145.61 19.19 0.496 0.707 147.10
Table 9: Quantitative evaluation of solving deblurring of different strength of degradation. We set $\eta_{1}=-0.4$ and $\eta_{2}=0$ for all of the tasks.
Kernel size DDNM MAS
PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ FID $\downarrow$
25 41.60 0.063 0.967 44.82 41.71 0.042 0.970 27.93
50 33.77 0.186 0.888 90.29 34.40 0.158 0.896 70.26
100 28.61 0.296 0.812 104.06 29.13 0.272 0.819 93.21
150 25.74 0.363 0.768 131.31 26.12 0.819 0.774 119.11
Figure 8: Super-resolution restoration over various strength of degradation. We set $\eta_{1}=-0.4$ and $\eta_{2}=0$ for all tasks. For sampling process, we set $\eta=0.6$.
G.4 Adding Gaussian noise
We present the results of solving 8×8$\times$8 × super-resolution under varying levels of Gaussian noise in Fig. 9. The visualizations demonstrate that MAS maintains strong restoration performance, even under high noise conditions.
Figure 9: MAS for solving super-resolution (8×8$\times$8 ×) with various strength of Gaussian noise.
G.5 Adding non-differentiable degragation
We present the results of solving 8×8$\times$8 × super-resolution with non-differentiable degradations, including JPEG compression and quantization, in Fig. 10.
Figure 10: Additional visualization of super-resolution with unknown noise.
Appendix H Licenses
FFHQ Dataset. We use the Flickr-Faces-HQ (FFHQ) dataset released by NVIDIA under the Creative Commons BY-NC-SA 4.0 license. The dataset is intended for non-commercial research purposes only. More details are available at: https://github.com/NVlabs/ffhq-dataset.
ImageNet Dataset. The ImageNet dataset is used under the terms of its academic research license. Access requires agreement to ImageNet’s data use policy, and redistribution is not permitted. More information is available at: https://image-net.org/download.